# Fate of the Unbound States in Near-infinitely Deep Potential Models

### 23 Apr 2026 v1

@@ III Non-Hermitian effect on unbound states
Figure 6: (Color online) The real part energies $E$ versus $V$ of the non-Hermitian deeper potential model. The colors denote the values of the IPR. The two blue solid lines denote the critical energies $E^{\rm NH2}_{c1}=2t-V$ and $E^{\rm NH2}_{c2}=-2t-V$, respectively. The energy domain $\left[E_{c2},E_{c1}\right]$ is a mixed region where unbound states and bound states coexist. Above $E_{c1}$ and below $E_{c2}$, there are bound states. The system size is $L=1597$ and $h=0.1$.
Figure 7: (Color online) Four wave functions of the non-Hermitian deeper potential model. (a) The unbound state at E=3.1524tE=-3.1524t; (b) The unbound state at E=0.7393tE=-0.7393t; (c) The bound state at $E=0.1940t$; (d) The bound state at $E=2.2617t$. The involved parameters are $h=0.1$, $V=t$, and $L=1597$.
Due to $\gamma^{B}_{\rm NH2}$ is a convex function that is linear in sections and has an integer slope, is can be concluded that the LE can be obtained when $\epsilon$ goes back to 0. Therefore, the LE $\gamma^{\rm NH2}(E)$ is
@@
From the above Eq. (28), we know that when $|E+V|>2t$, $\gamma^{\rm NH2}(E)>0$; while when $|E+V|<2t$, $\gamma^{\rm NH2}(E)=0$. The analytical results predict that the unbound states exist in the real energy domain 2tV<E<2tV-2t-V<E<2t-V and while the bound states will appear in the real energy domain $E>2t-V\bigcup E<-2t-V$. The expression of this solution is consistent with the result of Hermitian theory. However, it should be emphasized that for non-Hermitian models, the energy here refers to its real part. This will be fully reflected in the subsequent IPR calculations. For convenience, we mark $E^{\rm NH2}_{c1}=2t-V$ and $E^{\rm NH2}_{c2}=-2t-V$.
@@
In Fig. 6, we plot the real part energies $E$ (arranged in ascending order) as a function of the potential parameter $V$. The color denotes the magnitude of the IPR, and the blue lines mark the critical energies $E^{\rm NH2}_{c1}$ and $E^{\rm NH2}_{c2}$ obtained from analytical derivations. Similar to the Hermitian scenario, bound states still occupy the energy domain $E>E^{\rm NH2}_{c1}$. Differently, bound states emerge in the lower real part energy domain with $E<E^{\rm NH2}_{c2}$. In this regard, the distribution characteristics of the bound states are quite similar to that of the non-Hermitian Liu-Xia model. Furthermore, it is seen that the unbound states do not vanish due to non-Hermiticity, but remain present in the real part energy window $E^{\rm NH2}_{c2}<E<E^{\rm NH2}_{c1}$. Consistent with the non-Hermitian Liu-Xia model, the unbound states here still appear in the form of a mixed phase consisting of both unbound and bound states without a regular pattern. It is readily seen that the real part energies of unbound states are embedded in the real part energies of bound states as well. The emergent mixed phase once again indicates that non-Hermitian nature can fundamentally reconstruct the energy spectrum and localization structure of the system, forming the unconventional phenomenon.
@@
The unbound states, bound states, and mixed phase can all be reflected by the spatial distributions of the wave functions. Figures 7(a)-7(d) show the spatial distributions of the wave functions under four different real part energies. In Figure 7(a), the real part energy corresponding to the wave function lies below the critical energy $E^{\rm NH2}_{c2}$. It can be observed that its spatial distribution is localized, indicating a bound state. The wave functions in Figs. 5(b)-5(c) fall within the real part energy domain $E^{\rm NH2}_{c2}<E<E^{\rm NH2}_{c1}$. Concretely, the distribution in Fig. 5(b) is localized, corresponding to a bound state, while that in Fig. 5(c) is extended, corresponding to an unbound state. For fixed system parameters, the coexistence of both bound and unbound states in the energy spectrum characterizes the mixed phase. When the real part energy is larger than $E^{\rm NH2}_{c1}$, Fig. 7 shows that the wave function distribution is localized, corresponding to a bound state. These results clearly demonstrates that in the non-Hermitian deeper potential model, bound states and unbound states coexist within the energy domain $E^{\rm NH2}_{c2}<E<E^{\rm NH2}_{c1}$, which is consistent with the IPR results. This also indicates that the analytical solution derived based on Avila’s global theory can predict both the real part energy domain of unbound states and that of pure bound states.
@@ IV Summary
This paper focuses on the quasi-periodic nearly infinitely deep potential model, and systematically investigates the regulation of unbound states by potential well depth and non-Hermiticity via theoretical derivation and numerical simulation. We first construct a deeper potential well model based on the Liu-Xia model. Using Lyapunov exponent analysis, inverse participation ratio calculations and wave function characterization, we verify that deepening the potential well only compresses the energy range of bound states, while unbound states remain stable in specific energy domains. We then extend the study to non-Hermitian systems with on-site gain-loss effects. Numerical and analytical results show that non-Hermiticity does not eliminate unbound states, but leads to the formation of a mixed phase where bound and unbound states coexist. In addition, the analytical solution based on Avila’s global theory can accurately identify the mixed phase and pure bound states. This work reveals the modulation of extreme potentials and non-Hermiticity on unbound states, and provides new physical understanding and theoretical support for investigating bound-unbound characteristics in quantum systems.

### 27 Apr 2026 v2

@@ III Non-Hermitian effect on unbound states
Figure 6: (Color online) The real-part energies $E$ versus $V$ of the non-Hermitian deeper potential model. The colors denote the values of the IPR. The two blue solid lines denote the critical energies $E^{\rm NH2}_{c1}=2t-V$ and $E^{\rm NH2}_{c2}=-2t-V$, respectively. The energy domain $\left[E_{c2},E_{c1}\right]$ is a mixed region where unbound states and bound states coexist. Above $E_{c1}$ and below $E_{c2}$, there are bound states. The mixed interval follows the shifted analytical boundaries inherited from the deeper potential. The system size is $L=1597$ and $h=0.1$.
Figure 7: (Color online) Four wave functions of the non-Hermitian deeper potential model. (a) The bound state at E=3.1524tE=-3.1524t; (b) The bound state at E=0.7393tE=-0.7393t; (c) The unbound state at $E=0.1940t$; (d) The bound state at $E=2.2617t$. The parameters used are $h=0.1$, $V=t$, and $L=1597$.
Since $\gamma^{B}_{\rm NH2}$ is a convex function that is linear in sections and has an integer slope, it can be concluded that the LE can be obtained when $\epsilon$ goes back to 0. Therefore, the LE $\gamma^{\rm NH2}(E)$ is
@@
From the above Eq. (28), we know that when $|E+V|>2t$, $\gamma^{\rm NH2}(E)>0$; while when $|E+V|<2t$, $\gamma^{\rm NH2}(E)=0$. The analytical results predict that the unbound states exist in the real-energy domain 2tV<E<2tV-2t-V<E<2t-V, while the bound states appear in the real-energy domains E>2tVE>2t-V and E<2tVE<-2t-V. This expression is consistent with the result of the Hermitian theory. However, it should be emphasized that for non-Hermitian models, the energy here refers to its real part. This is fully reflected in the subsequent IPR calculations. For convenience, we mark $E^{\rm NH2}_{c1}=2t-V$ and $E^{\rm NH2}_{c2}=-2t-V$.
@@
In Fig. 6, we plot the real-part energies $E$ (arranged in ascending order) as a function of the potential parameter $V$. The color indicates the magnitude of the IPR, and the blue lines mark the critical energies $E^{\rm NH2}_{c1}$ and $E^{\rm NH2}_{c2}$ obtained from analytical derivations. Similar to the Hermitian case, bound states still occupy the energy domain $E>E^{\rm NH2}_{c1}$. Differently, bound states emerge in the lower real part energy domain with $E<E^{\rm NH2}_{c2}$. In this regard, the distribution characteristics of the bound states are quite similar to that of the non-Hermitian Liu-Xia model. Furthermore, it is evident that the unbound states do not vanish due to non-Hermiticity, but remain present in the real-part energy window $E^{\rm NH2}_{c2}<E<E^{\rm NH2}_{c1}$. Consistent with the non-Hermitian Liu-Xia model, the unbound states here still appear in the form of a mixed phase consisting of both unbound and bound states without a regular pattern. It is readily seen that the real-part energies of unbound states are embedded in the real-part energies of bound states as well. The reappearance of the mixed phase indicates that the non-Hermitian reconstruction of localization properties is robust against the deepening of the quasiperiodic potential.
@@
The unbound states, bound states, and mixed phase can all be reflected by the spatial distributions of the wave functions. Figures 7(a)-7(d) show the spatial distributions of the wave functions under four different real-part energies. In Fig. 7(a), the real-part energy corresponding to the wave function lies below the critical energy $E^{\rm NH2}_{c2}$. It can be observed that its spatial distribution is localized, indicating a bound state. The wave functions in Figs. 7(b)-7(c) fall within the real-part energy domain $E^{\rm NH2}_{c2}<E<E^{\rm NH2}_{c1}$. Concretely, the distribution in Fig. 7(b) is localized, corresponding to a bound state, while that in Fig. 7(c) is extended, corresponding to an unbound state. For fixed system parameters, the coexistence of both bound and unbound states in the energy spectrum characterizes the mixed phase. When the real-part energy is larger than $E^{\rm NH2}_{c1}$, Fig. 7(d) shows that the wave function distribution is localized, corresponding to a bound state. These results clearly demonstrate that in the non-Hermitian deeper potential model, bound states and unbound states coexist within the energy domain $E^{\rm NH2}_{c2}<E<E^{\rm NH2}_{c1}$, which is consistent with the IPR results. This also indicates that the analytical solution derived based on Avila’s global theory can predict both the real-part energy domain of unbound states and that of pure bound states.
@@ IV Summary
In summary, we have investigated the fate of unbound states in one-dimensional quasiperiodic models with near-infinitely deep potentials. For the deeper Hermitian model, numerical IPR calculations and wave-function profiles show that increasing the potential depth does not remove the unbound states. Instead, Lyapunov-exponent analysis based on Avila’s global theory gives the shifted unbound-state window 2tV<E<2tV-2t-V<E<2t-V, in agreement with the numerical phase diagram. Thus, the existence of unbound states is not determined simply by the absolute depth of the potential, but by the analytical energy window fixed by the Lyapunov exponent.
We have further extended both the Liu-Xia model and its deeper counterpart to non-Hermitian quasiperiodic potentials with gain and loss. In these cases, the relevant analytical boundaries are defined in terms of the real part of the complex eigenenergy. The Lyapunov exponent identifies the real-energy intervals in which unbound states can appear, while numerical IPR results reveal that these intervals become mixed bound-unbound phases rather than purely unbound regions. In contrast to conventional mobility-edge scenarios, localized and extended eigenstates coexist irregularly within the same real-energy window. Our results therefore show that non-Hermiticity does not destroy the unbound states in near-infinitely deep potentials, but reshapes their spectral organization and produces mixed localization structures. These findings provide a theoretical basis for further studies of bound-unbound transitions in extreme quasiperiodic and non-Hermitian systems.